### v1

@@ V-A Human-based Evaluation
In the domain of large model safety evaluation, several critical safety concerns are involved, including Toxicity, Ethics, and Bias. A central challenge in this context is defining the "safety" of the output content. Human-based evaluation holds a unique advantage in this regard due to its capacity for nuanced understanding and judgment of complex content [208]. Specifically, human evaluators are better equipped to interpret subtle contextual cues, tone, and sarcasm, enabling them to more accurately identify the potential meanings embedded in the content. This is particularly important for the detection of issues such as Toxicity and Bias. However, human evaluation is not without its challenges, primarily related to cost and consistency. It necessitates significant human and time resources [207], and in large-scale evaluations, as the volume of content increases, the associated costs also rise substantially. Moreover, due to variations in evaluators’ cultural backgrounds, knowledge bases, and personal values, human evaluation may lack consistency, particularly when assessing issues like Toxicity and Bias [74].
While automated evaluation offers advantages in terms of scalability and cost-effectiveness, it remains limited in its ability to fully comprehend complex contexts or diverse content. For example, models often struggle to grasp sarcasm, puns, and other intricate expressions, whereas human evaluators can provide a more comprehensive understanding in such cases. This limitation of automated methods is one of the reasons why many studies on large language model (LLM) safety continue to incorporate human evaluation components. For instance, Meadows et al. [15] utilized human evaluation to assess the alignment of LLM-generated content with local values, with a focus on evaluating the adherence of LLMs to Australian values and providing a framework for global regulatory bodies. Similarly, Alghamdi et al. [28] conducted a human-based evaluation of LLM credibility in Arabic-speaking contexts, finding that open-source models such as AceGPT 7B and Jais 13B struggled to achieve scores above 60% in benchmark tests. During their evaluation, they analyzed the cognitive differences in the perception of offensive content based on annotators’ gender, age, and cultural backgrounds. In a related study, Movva et al. [209] performed human evaluations where annotators assessed chatbot responses according to five safety standards—namely harm, bias, misinformation, political stance, and polarization—investigating the consistency between LLMs and human annotators in labeling conversational safety.
@@ V-B Rule-based Evaluation
Rule-based evaluation in the safety assessment of large language models typically involves the detection of potential risks or inappropriate content in model outputs by defining a set of explicit rules or guidelines. This approach relies heavily on predefined criteria to flag outputs that may contain sensitive information, inappropriate language, or other safety concerns. The process of rule-based evaluation generally follows these key steps:
  1. 1.
    Rule Definition: A comprehensive set of rules is established to clearly delineate which content is deemed unsafe, specifying the expressions or behaviors that are considered unacceptable.
  2. 2.
    Model Behavior Analysis: The model’s responses are compared against the predefined rules to identify potential violations or unsafe content.
  3. 3.
    Scoring Evaluation: Upon analyzing the model’s outputs, the content is evaluated and scored based on specific metrics to assess the overall safety performance of the model. Common metrics include accuracy, recall, and specific matching scores, such as ROUGE-L[90].
For example, Liu et al. [37] developed a set of evaluation rules designed to assess the copyright compliance of text generation. Specifically, they measured the similarity between generated content and copyrighted material using metrics like the Longest Common Substring (LCS) [226] and ROUGE-L scores[90]. These metrics help determine whether the model reproduces copyrighted text. In addition, the authors utilized multiple performance indicators, including LCS, ROUGE-L scores, and rejection rate. While LCS and ROUGE-L scores quantify the similarity of the generated content, the rejection rate reflects the model’s ability to effectively reject the generation of copyrighted content.
Similarly, Ye et al. [210] established a set of security rules encompassing three stages: input, execution, and output, to clearly define what content is considered unsafe. For each stage, multiple safety scenarios were outlined, with specific rules addressing issues such as the rejection of malicious queries, avoidance of incorrect tool selection, and filtering of harmful or erroneous information. The authors also quantitatively evaluated the model’s safety performance, utilizing metrics such as rejection rate, tool selection error rate, and the proportion of unsafe outputs.
While the rules-based evaluation method offers clear and interpretable criteria, its reliance on predefined rules presents limitations, particularly in detecting new, unknown risks or latent safety issues. As a result, rule-based evaluation is often combined with other evaluation methods, such as human evaluation and large model evaluation, to provide a more comprehensive and efficient assessment of the safety of large language models.
@@ V-C Model-based Evaluation
Model-based evaluation refers to the utilization of high-performance models to assess the safety of text generated by large language models (LLMs), such as GPT-4 [78], BERT [91], and reward models [216]. In this approach, models function as "evaluators" that interpret, analyze, and determine whether the generated content adheres to specific safety standards. Model-based evaluation offers substantial advantages in assessment capabilities [211, 212, 213]: in contrast to rule-based evaluation, it can handle complex semantic structures and context-dependent nuances, thereby identifying content with implicit biases or harmful elements. Furthermore, compared to human evaluation, model-based evaluation can process large volumes of data rapidly and efficiently, significantly reducing the need for human involvement and associated costs.
However, model-based evaluation is not without its limitations. Its evaluation process depends on sophisticated model architectures and hyperparameters, which can render the decision-making process opaque and difficult to interpret. Moreover, the training data used to develop the evaluation models may carry inherent biases [215], which could influence the evaluation results, leading to potential inconsistencies, especially when assessing sensitive or controversial topics.
V-C1 Closed-source Model
Closed-source models, as evaluators, offer high performance and versatility. Models like GPT-4 and InstructGPT, which are trained on extensive datasets, exhibit remarkable capabilities in understanding language and analyzing complex contexts. These models excel in tasks such as sentiment analysis, semantic interpretation, and intent detection, making them highly effective for evaluating content safety and quality. It cannot be denied that GPT-4-as-a-judge is the most popular evaluation method in the field of jailbreak attacks and has been proven to have strong correlation with human judgment. In addition, there are other ways to use closed source models to assist in evaluation, for instance:
Cai et al. [177] introduced three novel evaluation metrics—Safeguard Violation (SV), Informativeness (I), and Relative Truthfulness (RT)—to more precisely assess model jailbreak outcomes. The evaluation methodology leverages GPT-4 as the evaluator, utilizing predefined prompt templates to assess these three metrics. Several other studies [49, 217, 186, 21] have similarly employed GPT-4 as an evaluation tool. Sun et al. [17] proposed a benchmarking framework for assessing the safety of Chinese LLMs, with InstructGPT serving as the evaluator. Through dynamic prompt templates, InstructGPT is tasked with evaluating the safety of responses across various scenarios. The safety response rate for each scenario is calculated and aggregated to produce an overall safety ranking.
Despite their advantages, closed-source models present significant challenges regarding transparency and interpretability. The lack of public disclosure regarding their internal architectures and training data makes it difficult to offer clear explanations for the evaluation outcomes. This opacity can undermine trust, particularly in the context of sensitive content evaluation. Additionally, closed-source models have limited customization capabilities, which restrict their adaptability to specific evaluation needs, thus hindering their applicability in highly specialized contexts. While these models are well-suited for general evaluation tasks, their limitations in transparency and customization must be carefully considered when determining their use in particular scenarios.
V-C2 Open-source Model
@@
Several researchers have released open-source models to support safety evaluation efforts. These models are often fine-tuned on specific datasets to approach the performance of state-of-the-art closed-source models, and they can be deployed locally, reducing evaluation costs while improving reproducibility.
@@
Llama Guard [137] is a safety evaluation model fine-tuned from the Llama2-7B model. It classifies responses based on user prompts and model outputs, and for responses labeled as unsafe, it also provides the violated safety policy. Meta researchers continue to update this series, with versions including Llama Guard 2 8B [219], Llama Guard 3 1B [220], and Llama Guard 3 8B [221].
ShieldLM [222] fine-tuned Qwen-14B-Chat [223] on a bilingual query-response dataset. Unlike most works that label responses as simply safe/unsafe, ShieldLM introduces a new label, "controversial." Under strict rules, controversial responses are marked as unsafe, while under loose rules, they are marked as safe. The model outputs both classification results and reasoning analyses.
MD-Judge [47], inspired by Llama Guard [137], fine-tuned the Mistral-7B model [224] for safety evaluation. Specifically, the researchers incorporated attack-augmented data in the fine-tuning process to further enhance the model’s capabilities. The model outputs binary classification labels of safe/unsafe and the category of the unsafe response.
AegisSafetyExperts [218] is unique in that its goal is to build a collection of LLM-based safety experts. The researchers fine-tuned three models using their own constructed dataset. First, based on Llama Guard [137], they fine-tuned two versions using the proposed taxonomy and safety policies—Llama Guard Defensive and Llama Guard Permissive. The former maps safety categories that “Need Cation” to unsafe, while the latter maps them to safe. Additionally, the researchers fine-tuned the NeMo43B model using their taxonomy in the prompt, resulting in the NeMo43B-Defensive model for prompt classification.
WildGuard [51] also fine-tuned Mistral-7B [224] to detect malicious intent in user prompts, assess safety risks in model responses, and determine the model’s refusal rate. WildGuard’s training data includes synthetic adversarial, synthetic vanilla, and real-world user-LLM interaction data to maximize task coverage, diversity, and balance.
ShieldGemma [225] is designed specifically for safety content moderation. The researchers supervised fine-tuning of Gemma2 (2B, 9B, and 27B parameters), requiring outputs to include binary classification results and reasoning. Experiment results show that ShieldGemma outperforms both Llama Guard[137] and WildGuard[51] .
JailJudge Guard [194] is the latest research, fine-tuning the Llama2-7B model with instruction tuning. The model is trained to output detailed reasoning and a fine-grained evaluation score (a jailbreak score from 1 to 10). To further showcase JailJudge Guard’s capabilities, the researchers developed an attacker-agnostic attack enhancer and a system-level jailbreak defense method.
@@ VI-A Unified Evaluation
Currently, the research community has not reached a consensus on how to conduct specific safety evaluations for LLMs. There are differences in the specific implementation of evaluation methods and benchmarks used by different researchers. Especially for widely used key metrics such as attack success rate, there is no clear and specific measurement method. Such inconsistency makes it difficult to compare work across researchers and even raises questions about the reliability and accuracy of evaluation results[184]. Furthermore, as the application areas of LLMs continue to expand, the safety evaluation requirements vary across different fields, exacerbating these inconsistencies. Therefore, we believe it is urgent to develop a universal evaluation system that can support various task types, precisely calculate metrics, and adapt flexibly to the evaluation needs of different models and domains. Although it may seem like an unlikely goal to complete, future research should strive to approximate this ultimate objective by establishing standardized evaluation procedures, unified and comprehensive metrics, and a multi-layered evaluation framework.

### v2

@@ V-A Human-based Evaluation
Since the birth of the field of machine learning, human judgment has been regarded as the gold standard. In the safety evaluation of LLMs, central concerns include toxicity, ethics, and bias. A persistent challenge in this area is defining what “safety” means in relation to model outputs. Human-based evaluation holds a unique advantage here, as it allows for nuanced interpretation and context-sensitive judgment that automated approaches struggle to replicate [189]. Human evaluators can recognize subtle cues such as tone, sarcasm, or implicit meanings, making them particularly well-suited to detecting toxicity and bias. Nevertheless, human evaluation faces significant challenges, especially in terms of cost and consistency. It necessitates significant human and time resources [190], and in large-scale evaluations, as the volume of content increases, the associated costs also rise substantially. For instance, Yuan et al. [134] compensate annotators at a rate of \$27.5 per hour according to the standard hourly income for U.S. annotators. Moreover, variations in cultural background, knowledge, and personal values can reduce consistency across annotators, particularly in sensitive domains such as toxicity and bias [109].
Despite these limitations, human evaluation remains indispensable because automated methods are still constrained in their ability to grasp complex or culturally embedded content. Models often fail to interpret sarcasm, puns, or layered meanings, whereas human annotators can provide a more comprehensive understanding. This explains why many LLM safety studies continue to integrate human judgment into their evaluation pipelines. For instance, Meadows et al. [36] relied on human evaluation to examine whether LLM outputs aligned with Australian values, offering insights for global regulatory bodies. Similarly, Movva et al. [87] used annotators to evaluate chatbot responses across five safety dimensions—harm, bias, misinformation, political stance, and polarization—analyzing the agreement between human and model-based assessments. Human judgment is also a key reference point in jailbreak attack research, where it serves as the benchmark for evaluating both the success of jailbreaks and the reliability of automated evaluators. For example, Souly et al. [63] showed strong alignment between their proposed method and human judgment by comparing bias levels and mean absolute error (MAE) against human evaluation results.
@@ V-B Rule-based Evaluation
Rule-based evaluation in the safety assessment of LLMs typically involves the detection of potential risks or inappropriate content in model outputs by defining a set of explicit rules or guidelines. Such methods rely on well-specified criteria to flag responses that may contain sensitive information, harmful language, or other safety concerns. The typical process involves defining rule sets, applying them to model outputs, and categorizing responses accordingly.
Among these methods, string matching remains the most widely adopted, particularly in jailbreak attack research. In this setting, evaluators probe models with harmful queries and determine the success of the jailbreak attempt by checking for the presence of “rejection strings” in the output (e.g., “I am sorry” or “As an AI”). Each study typically constructs its own phrase set, often based on heuristic observations from empirical experiments. String matching is favored for its interpretability and cost-effectiveness. However, it suffers from significant drawbacks. Most notably, it exhibits a high false positive rate and tends to overestimate unsafe behavior, with low correlation to human judgment. This occurs because rejection phrase lists cannot exhaustively capture all valid refusal expressions. For instance, the list employed by Zou et al. [8] excludes common phrases such as “not legal” and “not ethical”,which may lead to misclassification of safe outputs as harmful. Furthermore, jailbreak prompts often embed benign content to obscure malicious intent. In such cases, model responses may be irrelevant or unhelpful to the attack but are nevertheless flagged as unsafe under string-matching rules. Despite these limitations, string matching remains widely used, largely due to its negligible implementation cost and the need for comparability with prior studies that have adopted the same methodology [191, 133]. In addition, string matching has been observed to be combined with other security evaluation methods. For example, Liu et al. [88] combined string matching with text classifiers, while Zeng et al. [89] used string matching as a preliminary filter before conducting a more precise evaluation with GPT-4.
Beyond string matching, rule-based evaluation can also integrate traditional machine learning metrics to assess more specialized risks. For example, Liu et al. [60] proposed a framework to evaluate copyright compliance in text generation. Their method combined multiple rule-based indicators, including similarity metrics such as the Longest Common Substring (LCS) [192] and ROUGE-L scores [125], to quantify the overlap between generated text and copyrighted material. They further incorporated rejection rates as an additional measure of the model’s capacity to deny requests for infringing outputs. This integration of rule-based and metric-based evaluation provides a more comprehensive perspective than string matching alone.
In summary, rule-based evaluation offers clear, interpretable, and cost-efficient criteria for identifying unsafe or noncompliant content in LLM outputs. Nonetheless, its dependence on predefined rules makes it inherently limited in detecting novel risks or subtle safety issues. As such, rule-based methods are most valuable when employed as baseline or complementary tools, often in combination with human or model-based evaluation, to ensure more robust and scalable safety assessments.
@@ V-C Model-based Evaluation
Model-based evaluation refers to the use of high-performance models to assess the safety of text generated by LLMs, such as GPT-4 [111], BERT [193], and reward models [194]. In this paradigm, models function as “evaluators” that interpret, analyze, and determine whether generated content adheres to predefined safety standards. Compared with rule-based evaluation, model-based methods can capture complex semantic structures and context-dependent nuances, thereby identifying subtle biases and harmful contents [195, 196, 197]. Moreover, in contrast to human evaluation, model-based approaches scale efficiently to large volumes of data, reducing human labor and associated costs. Broadly, model-based evaluation techniques can be divided into three categories: (i) specialized text classifiers, (ii) closed-source chat models, and (iii) open-source chat models.
V-C1 Text Classifiers
Text classifiers generate structured outputs such as labels or confidence scores, which makes them particularly well-suited for safety evaluation tasks.
Several companies provide proprietary content moderation services for detecting harmful text. Widely used tools include OpenAI’s Moderation Endpoint [90], Microsoft’s Azure AI Content Safety API [92], and Google’s Perspective API [91]. For instance, OpenAI’s Moderation Endpoint is designed to determine whether content complies with its usage policies. It evaluates 11 categories—including hate, harassment, self-harm, sexual content, and violence—returning both binary labels and confidence scores between 0 and 1. Google’s Perspective API adopts a similar approach, focusing on “toxicity,” which refers to rude, disrespectful, or otherwise inappropriate text, but leaves the choice of decision threshold to users. While such closed-source classifiers are convenient, their utility for research purposes is limited. They primarily focus on surface-level harmful content, which can result in false positives in scenarios where unsafe behavior occurs implicitly (e.g., jailbreak demonstrations without overtly harmful text) [93].
In contrast, open-source classifiers are more widely used in academic evaluations. These models are typically fine-tuned from established sequence classification architectures such as BERT, RoBERTa, or DeBERTa. With parameter counts often below one billion, they are significantly smaller than contemporary open-source chat models. For example, Wang et al. [67] fine-tuned a Longformer [198] for harmful response detection, while Yu et al. [93] fine-tuned a RoBERTa-large model [130] for evaluation tasks. Nevertheless, the effectiveness of such classifiers heavily depends on the quality of their training datasets, and their performance often degrades on out-of-distribution data.
V-C2 Closed-source Chat Model
Closed-source chat models, when employed as evaluators, provide both high performance and versatility. Models such as GPT-4 and InstructGPT, trained on massive datasets, demonstrate strong abilities in contextual understanding, semantic interpretation, and intent recognition. As a result, they are increasingly adopted as evaluators for assessing the safety and quality of model outputs. In particular, GPT-4-as-a-judge has become the most widely used evaluation method in this domain, showing strong correlation with human judgments.
The implementation details of the evaluation process are all based on prompt engineering. By designing an evaluation prompt word, the model is prompted to evaluate the input-output pair according to the evaluation rules. For example, the GPT-4 Judge prompt [94] uses a Likert-style prompt to assess the harmfulness of a response. The model is informed of OpenAI’s relevant safety policies as a reference for the evaluation process. PAIR prompt [95] assesses the degree to which the model’s response violates ethical, legal, or safety guidelines, outputting a score out of 10. Similarly, PICT [96] assesses whether the victim model’s response contains harmful information related to a given category of prohibited behavior.
Despite their advantages, closed-source models face challenges in reliability. Prior studies highlight issues such as limited reproducibility, sensitivity to prompt wording, and a lack of robustness [197, 199, 200]. These shortcomings undermine consistency across evaluations and raise concerns regarding reproducibility and transparency in academic or regulatory contexts. While closed-source models remain powerful and cost-efficient alternatives to human evaluation, further research is needed to enhance their robustness and reliability.
V-C3 Open-source Chat Model
@@
Several researchers have released open-source chat models to support safety evaluation efforts. These models are often fine-tuned on specific datasets to approach the performance of state-of-the-art closed-source models, and they can be deployed locally, reducing evaluation costs while improving reproducibility. Existing works can be broadly divided into two categories: models focusing on binary results with classification labels, and models that emphasize diversified labels and detailed reasoning.
@@
Binary results and classification labels. Llama Guard [97] is a safety evaluation model fine-tuned from the Llama2-7B model. It classifies responses based on user prompts and model outputs, and for responses labeled as unsafe, it also provides the violated safety policy. Meta researchers continue to update this series, with versions including Llama Guard 2 8B [98], Llama Guard 3 1B [99], and Llama Guard 3 8B [100]. Similarly, MD-Judge [72], inspired by Llama Guard, fine-tuned the Mistral-7B model [201] for safety evaluation. Specifically, the researchers incorporated attack-augmented data in the fine-tuning process to further enhance the model’s capabilities. The model outputs binary classification labels of safe/unsafe and the category of the unsafe response.
Diversified results and reasoning. Other works go beyond binary classification and focus on providing richer evaluation outcomes. ShieldLM [101], fine-tuned from Qwen-14B-Chat [202], introduces a new label “controversial.” Under strict rules, controversial responses are marked as unsafe, while under loose rules, they are marked as safe. The model outputs both classification results and reasoning analyses. AegisSafetyExperts [71] is unique in that its goal is to build a collection of LLM-based safety experts. The researchers finetuned three models using their own constructed dataset. First, based on Llama Guard, they fine-tuned two versions using the proposed taxonomy and safety policies — Llama Guard Defensive and Llama Guard Permissive. The former maps safety categories that “Need Cation” to unsafe, while the latter maps them to safe. Additionally, the researchers fine-tuned the NeMo43B model for prompt classification. WildGuard [74] also fine-tuned Mistral-7B to detect malicious intent in user prompts, assess safety risks in model responses, and determine the model’s refusal rate. Its training data includes synthetic adversarial, synthetic vanilla, and real-world user-LLM interaction data to maximize task coverage, diversity, and balance. ShieldGemma [102] is designed specifically for safety content moderation. The researchers supervised fine-tuning of Gemma2 (2B, 9B, and 27B parameters), requiring outputs to include binary classification results and reasoning. Experiment results show that ShieldGemma outperforms both Llama Guard and WildGuard. Finally, JailJudge Guard [69] represents the latest research in this line. By fine-tuning the Llama2-7B model with instruction tuning, it produces detailed reasoning together with a fine-grained evaluation score (a jailbreak score from 1 to 10).
Overall, open-source chat models present a flexible alternative for automated safety evaluation. The primary differences among existing models lie in the composition of training data and the form of evaluation outputs. However, their deployment requires access to large-scale, high-quality labeled datasets and significant computational resources for fine-tuning. Moreover, their generalization to out-of-distribution data remains an open challenge, raising questions about their reliability in real-world applications.
V-D Evaluation Frameworks
Evaluation frameworks integrate the entire evaluation pipeline, enabling convenient end-to-end assessment and simplifying the process. These frameworks take evaluation datasets as input and produce evaluation results as output. In the intermediate process, in addition to interacting with target models through interfaces (such as HF, VLLM, or OpenAI) to obtain outputs and conduct evaluation, some frameworks also include components for attack augmentation on data or defense enhancement on models. These are designed to test the robustness of target models against attacks and the effectiveness of defense methods. As a result, most evaluation frameworks are typically composed of four parts, or a subset of them: Attacks, Defenses, Judges, and Interfaces.
As shown in Table III, we summarize several representative evaluation frameworks and provide a brief comparison. JailbreakEval [17] is a toolkit dedicated to evaluating jailbreak attempts. It does not include a preliminary data processing stage but instead focuses only on evaluating results. Its input is “question–answer” pairs, and it provides a set of out-of-the-box evaluators. In contrast, EasyJailbreak [103] is designed specifically for jailbreak attacks and consists of four components: Selector, Mutator, Constraint, and Evaluator. The process has three stages: in the preparation stage, users provide malicious prompts, prompt templates, or other configurations; in the attack stage, EasyJailbreak iteratively attacks the target model and updates the prompts while evaluating the outcome of each attack; in the final stage, the system aggregates results and generates a report for the user. Moving beyond general jailbreak-specific settings, WalledEval [104] introduces innovation by supporting a wide range of safety benchmarks, with more than 35 benchmarks built in. These cover multilingual safety, exaggerated safety, prompt injection, and red-teaming datasets. Finally, HarmBench [31] is a standardized framework for automated red-teaming. It emphasizes three qualities of automated red-team evaluation: breadth, comparability, and robustness. Based on these principles, the pipeline begins by converting a diverse set of behaviors into test cases to probe target models, and then applies multiple classifiers to evaluate the responses and calculate the attack success rate.
Table III: Comparison of Different Evaluation Frameworks.
Frameworks Attacks Defenses Judges Interfaces
JailbrakEval [17] - - 22 -
EasyJailbreak [103] 12 - 7 4
WalledEval [104] 9 - 6 6
HarmBench [31] 18 - 3 5
JailbreakBench [30] 4 5 6 3
AISafetyLab [105] 13 16 7 3
PandaGuard [106] 19 12 4 7
The above frameworks do not consider defenses. In contrast, JailbreakBench [30] incorporates several defense methods compared with EasyJailbreak. This toolkit allows users to configure threat models, system prompts, chat templates, and scoring functions, and it maintains a public leaderboard that compares the performance of state-of-the-art jailbreak attacks and defenses. Building on a broader perspective, AISafetyLab [105] is organized into three core modules — attack, defense, and evaluation — along with four shared auxiliary modules: models, dataset, utils, and logging. It provides researchers with a highly extensible framework. Further, PandasGuard [106] unifies attackers, defenders, target models, and judges into a modular system. Its main contribution is conducting a large-scale benchmark of about 3B tokens across 49 models.
In terms of evaluators, these frameworks generally adopt rule-based string matching and model-based evaluation methods. There is no clear preference between closed-source and open-source chat models, though text classifiers are used less frequently. While existing frameworks continue to expand in attacks, defenses, judges, and interfaces, there is still no consensus on evaluation methods. The divergence in evaluation method choices further highlights the lack of standardized practices in the field. This inconsistency not only weakens the comparability of results across frameworks but also limits the accumulation of research progress. Overall, current evaluation frameworks focus more on integrating and balancing existing methods than on advancing evaluation methodology. For higher-level evaluation, it is crucial to develop mechanisms that combine efficiency, accuracy, and robustness, and to establish standardized benchmarks and evaluation protocols.
@@ VI-A Unified Evaluation
Despite growing efforts in safety evaluation of LLMs, the research community has yet to reach consensus on standardized procedures and methodologies for conducting evaluations. Prior work has highlighted that that LLM safety evaluation lacks robustness due to noise stemming from factors such as small datasets, methodological inconsistencies, and unreliable evaluation setups [203]. There are differences in the specific implementation of evaluation methods and benchmarks used by different researchers. Especially for widely used key metrics such as attack success rate, there is no clear and specific measurement method. Such inconsistency makes it difficult to compare work across researchers and even raises concerns about reproducibility and reliability [63]. Furthermore, as LLMs are increasingly applied in diverse domains, ranging from education to healthcare and law, domain-specific safety requirements further exacerbate these inconsistencies. Therefore, we argue that future research should aim to establish a more unified evaluation paradigm. While achieving complete unification may be difficult, progress can be made by (i) developing standardized evaluation procedures and reporting protocols, (ii) defining unified and comprehensive metrics with clear computational formulations, and (iii) designing a multi-layered, modular evaluation framework that can flexibly adapt to different models, domains, and application scenarios. Such efforts would not only improve the comparability and reproducibility of safety evaluations but also foster a shared foundation for cumulative progress in this rapidly evolving field.